OpenStudy (anonymous):

Can someone Check my work for a differential equation work please? It ask to prove that the function is a solution to the differential equation. Give me a min to use the function builder to input the formulas please.

OpenStudy (anonymous):

\[\frac{dy}{dx}=ay\] assume a is a coefficient

OpenStudy (anonymous):

\[\frac{1}{ay}dy=dx\]

OpenStudy (anonymous):

\[\int\limits(\frac{1}{ay})dy=\int\limits dx\]

OpenStudy (anonymous):

\[\log(ay)=x+c\]

OpenStudy (anonymous):

\[y=\frac{e^{x+c}}{a} \]

OpenStudy (anonymous):

\[\lim_{y \rightarrow o}y=\lim_{x \rightarrow 0}\frac{e^{x+c}}{a}\]

OpenStudy (anonymous):

0=0

OpenStudy (anonymous):

So that proves that it's the solution right?

OpenStudy (turingtest):

no, because \[\lim_{x \rightarrow 0}\frac{e^{x+c}}{a}=\frac{e^c}a\]and we don't know what c is, and even if c was 0, that would give\[\lim_{x \rightarrow 0}\frac{e^{x+c}}{a}=\frac{e^0}a=\frac1a\]

OpenStudy (turingtest):

what is the function that you are trying to prove is a solution?

OpenStudy (turingtest):

I may not be online by the time you reply, but if you have a function for y, I suggest you simply plug it in to the differential equation above and make sure it comes out the same on both sides.

OpenStudy (turingtest):

that would prove it as a solution to the DE

OpenStudy (anonymous):

What makes this a pain is that I don't have the book so I'm using pictures that a class mate took for me. Let me post the picture see if you have any better luck understanding it than I do. It's problem number 1 on this page i'm about to upload.

OpenStudy (anonymous):

oh crap I didn't see the functions on the right of the diff eq

OpenStudy (loser66):

\[\frac{dy}{dx}= ay \] solution is \(y = e^{ax}\) right?